OpenStudy (anonymous):

A stone is dropped onto a lake, creating a circular ripple that travels outward at a speed of 50in=sec. Find the rate at which the area within the circular ripple is increasing after 1sec and 3sec

OpenStudy (anonymous):

So this is a typical related rates problem with an increasing circle..

OpenStudy (anonymous):

First write the formula for the Area of the circle: A = pi*r^2

OpenStudy (anonymous):

dA/dt is the rate at which the circle is increasing, which is also the derivative: dA/dt = 2*pi*r*(dr/dt)

OpenStudy (anonymous):

hmm actually what do they mean by is traveling outward? Does that mean the radius is increasing?

OpenStudy (anonymous):

correct i believe that is exactly how it is written, I was hoping someone might know what they meant.

OpenStudy (anonymous):

Oh i think it is the radius.. let us just work it out and see for ourselves..

OpenStudy (anonymous):

So: dA/dt = 2*pi*r*(dr/dt) the radius is increasing at the rate of 50 in/sec So in one second, the radius increased from 0 to 50in. So: dA/dt = 2*pi*50*50 = = 5000pi

OpenStudy (anonymous):

In 3 seconds radius increased from 0 to 150in. So: dA/dt = 2*pi*150*50 = = 15000 pi check the answers..

OpenStudy (anonymous):

ok that s what I was thinking but I didnt want to be wrong when I have a test in 45 min. thank you very much.